OpenStudy (anonymous):

Solve: negative 1 over 2 x + 2 = –x + 7

OpenStudy (anonymous):

Is the equation like this? \[-\frac{ 1 }{ 2x+2}=-x+7\]

OpenStudy (anonymous):